Calibration status update

Speaker : Gene Van Buren ( BNL )


Talk time : 12:15, Duration : 00:15
Run 13 pp510:
  • Initial non-zero (and tilted) BeamLine in place
  • Preliminary TPC SpaceCharge & GridLeak (using Run 12 pp510 final result) in place
    • initial calibration using Run 13 data ongoing
  • Next steps in the calibration: TPC T0 & Twist (no one working on this)
  • Some hiccups in the automatic TPC drift velocity calibration between March 21-27, but tuning the lasers seems to have helped since then
  • TPC group focused on bringing GMT software to readiness to use in calibrations (alignment, SpaceCharge & GridLeak, possibly drift velocity)
  • HFT & TPC groups interested in a low luminosity fill (in case cosmics are insufficient and/or AuAu15 does not happen)
    • seeking additional interest/input, and will post the full known list of requirements for the fill within ~1 day

Run 12 pp510:
  • BeamLine calibration done
  • dE/dx calibration:
    • production processed (with 2.5M events)
    • need to be careful about ZDC-based luminosity correction, i.e. days 85 & 86, as was the case for SC&GL calibration)
    • open question of whether -0.5 sigma shift is seen here too as it was in Run 12 pp200
      • this will be a key factor in how long this will take to complete! 
  • BTOF/VPD check of their calibration ongoing
    • hampered by VPD failures (day 87, as was an issue for SC&GL calibration too)
    • expecting to finish by the end of this week
  • TPC hit errors to be done last (~1 day)

-Gene